# Revealing the nature of $\Omega_{c}$-like states from pentaquark perspective

### v2

@@ Abstract
We systematically study the electromagnetic properties of controversial states whose internal structure is not elucidated and we try to offer a different point of view to unravel the internal structure of these states. Inspired by the $\Omega_{c}$ states observed by the LHCb Collaboration, we study the electromagnetic properties of the $\Omega_{c}$-like states as the compact diquark-diquark-antiquark pentaquarks with both $J^{P}=\frac{1}{2}^{-}$ and $J^{P}=\frac{3}{2}^{-}$ in the context of the QCD light-cone sum rule model. From the obtained numerical results, we conclude that the magnetic dipole moments of the $\Omega_{c}$-like states can reflect their inner structures, which can be used to distinguish their spin-parity quantum numbers. Measuring the magnetic moment of the $\Omega_{c}$-like states in future experimental facilities can be very helpful for understanding the internal organization and identifying the quantum numbers of these states.
Electromagnetic form factors, diquark-diquark-antiquark picture, QCD light-cone sum rules, magnetic dipole moments
@@ I Motivation
$M_{\Omega_{c}(3000)}=3000.4\pm 0.2\pm 0.1\mbox{ MeV}\,,\,\,\,\,\Gamma_{\Omega_{c}(3000)}=4.5\pm 0.6\pm 0.3\mbox{ MeV}\,,$
$M_{\Omega_{c}(3050)}=3050.2\pm 0.1\pm 0.1\mbox{ MeV},\,\,\,\,\Gamma_{\Omega_{c}(3050)}=0.8\pm 0.2\pm 0.1\mbox{ MeV}\,,\,\,$
$M_{\Omega_{c}(3066)}=3065.6\pm 0.1\pm 0.3\mbox{ MeV}\,,\,\,\,\,\Gamma_{\Omega_{c}(3066)}=3.5\pm 0.4\pm 0.2\mbox{ MeV}\,,$
$M_{\Omega_{c}(3090)}=3090.2\pm 0.3\pm 0.5\mbox{ MeV}\,,\,\,\,\,\Gamma_{\Omega_{c}(3090)}=8.7\pm 1.0\pm 0.8\mbox{ MeV}\,,$
$M_{\Omega_{c}(3119)}=3119.1\pm 0.3\pm 0.9\mbox{ MeV}\,,\,\,\,\,\Gamma_{\Omega_{c}(3119)}=1.1\pm 0.8\pm 0.4\mbox{ MeV}\,.$ (1)
@@
Later, the former four states $\Omega_{c}(3000)$, $\Omega_{c}(3050)$, $\Omega_{c}(3066)$ and $\Omega_{c}(3090)$ were confirmed by the LHCb and Belle Collaborations [31, 32]. The discovery of the LHCb collaboration has led to a new experimental situation, which requires a more detailed study of heavy baryons and their properties. The discovery of these states, although their quantum numbers have not been determined, could provide new insights into QCD and its complex behavior. This could lead to a deeper understanding of the underlying properties of QCD. However, our information about their properties is still not sufficient and further suggestions for experimental exploration of $\Omega_{c}$-like states should be addressed. The experimental discoveries were followed by various theoretical studies, investigating them in the conventional baryon, molecular, and compact pentaquark states to shed light on their exact nature and quantum numbers (for details see the Refs. [33, 34, 35, 36, 37]).
@@
The literature review indicates that the majority of research has concentrated on computing the spectroscopic and decay parameters of these states. However, it is evident that relying exclusively on spectroscopic and decay parameters is insufficient for resolving the disputed nature of these states. Therefore, to determine the precise internal configurations of these states, additional studies such as obtaining electromagnetic multipole moments, radiative decays, and weak decays are required. To unveil the nature and internal structures of unconventional states, the physical quantities associated with the electromagnetic properties of hadrons are valuable parameters. The charge and magnetism distributions inside the hadrons can be studied using the electromagnetic form factors (FFs) and the resulting multipole moments. This information can be used to determine the distribution of quark-antiquark pairs (both valence and sea-quarks) and gluons within the hadron volume. It also provides valuable insights into the geometric forms, electric radii, and magnetic radii of the hadrons. With this motivation, in this work, we will derive the electromagnetic features of $\Omega_{c}$-like states using the method of QCD light-cone sum rules [38, 39, 40]. Throughout the analysis, these states are considered in the compact $[ss][qc][\bar{q}]$, $[sq][sc][\bar{q}]$ and $[qs][sc][\bar{q}]$ diquark-diquark-antiquark configurations, and are assumed to have both $J^{P}=\frac{1}{2}^{-}$ and $J^{P}=\frac{3}{2}^{-}$ quantum numbers. There are several studies in the literature that have extracted the electromagnetic features of the singly-heavy nonconventional hadron states [41, 42, 43, 44, 45, 46, 47, 48].
@@
The article is structured as follows: we derive the QCD light-cone sum rules for the electromagnetic features of the $\Omega_{c}$-like states in Sect. II. The numerical results of the magnetic dipole moments and discussions are presented in Sect. III. Sect. IV is reserved for our summary and concluding remarks.
II Electromagnetic form factors from QCD light-cone sum rules
@@
The study of the electromagnetic form factors of the $\Omega_{c}$-like states ($\frac{1}{2}^{-}\rightarrow\Omega_{c}$, $\frac{3}{2}^{-}\rightarrow\Omega_{c}^{*}$) in QCD light cone sum rules begins with the writing of the correlation function in the weak electromagnetic background field ($\gamma$) as follows
@@
where $T$ is the time ordered product; and the $J^{\Omega_{c}}(x)$ and $J_{\mu}^{\Omega_{c}^{*}}(x)$ are the interpolating currents for the $\Omega_{c}$-like states. The explicit form of the $J^{\Omega_{c}}(x)$ and $J_{\mu}^{\Omega_{c}^{*}}(x)$ are written as follows
$J^{\Omega_{c}}(x)$ $=\varepsilon^{abc}\varepsilon^{ade}\varepsilon^{bfg}\big{[}q^{d^{T}}_{1}(x)C\gamma_{\mu}q_{2}^{e}(x)\,q_{3}^{f^{T}}(x)C\gamma^{\mu}c^{g}(x)\,C\bar{q}^{c^{T}}(x)\big{]},$ (4)
$J_{\mu}^{\Omega_{c}^{*}}(x)$ $=\varepsilon^{abc}\varepsilon^{ade}\varepsilon^{bfg}\big{[}q^{d^{T}}_{1}(x)C\gamma_{\mu}q_{2}^{e}(x)\,q_{3}^{f^{T}}(x)C\gamma_{5}c^{g}(x)\,C\bar{q}^{c^{T}}(x)\big{]},$ (5)
where $q_{1}$, $q_{2}$, $q_{3}$ are u-, d- and/or s-quark, $q$ is u- or d-quark, the $C$ is the charge conjugation operator and $a$, $b$… are color indices. It should be noted here that the mass values obtained with these interpolating currents are compatible with the mass values of the $\Omega_{c}(3050)$, $\Omega_{c}(3066)$, $\Omega_{c}(3090)$, and $\Omega_{c}(3119)$ states [49, 50].
@@
where $F_{M}(0)$ and $G_{M}(0)$ stand for magnetic form factor of the spin-$\frac{1}{2}^{-}$ and spin-$\frac{3}{2}^{-}$ $\Omega_{c}$-like states, respectively. Explicit forms of these terms are given as,
@@
$+\rm{Tr}\Big{[}\gamma^{\mu}S_{c}^{gg^{\prime}}(x)\gamma^{\nu}\widetilde{S}_{q_{2}q_{3}}^{ef^{\prime}}(x)\gamma_{\mu}S_{q_{1}q_{2}}^{de^{\prime}}(x)\gamma_{\mu}\widetilde{S}_{q_{3}q_{1}}^{fd^{\prime}}(x)\Big{]}\bigg{\}}\widetilde{S}_{q}^{c^{\prime}c}(-x)|0\rangle_{\gamma},$ (20)
@@
$+\rm{Tr}\Big{[}\gamma_{5}S_{c}^{gg^{\prime}}(x)\gamma_{5}\widetilde{S}_{q_{2}q_{3}}^{ef^{\prime}}(x)\gamma_{\mu}S_{q_{1}q_{2}}^{de^{\prime}}(x)\gamma_{\mu}\widetilde{S}_{q_{3}q_{1}}^{fd^{\prime}}(x)\Big{]}\bigg{\}}\widetilde{S}_{q}^{c^{\prime}c}(-x)|0\rangle_{\gamma},$ (21)
@@
where one of the propagators for either the light or heavy quark interacts with the photon at a short distance, and the surviving four propagators are assumed to be free.
In the case of the non-perturbative contributions, the following equation is to be used
@@
After completing the procedures described above, we have finished analytical calculations of the magnetic dipole moments for the relevant states. For simplicity, analytical calculations for the spin-1/2 $\Omega_{c}$ states are presented as an example in the appendix. The results for the magnetic dipole moments of $\Omega_{c}$-like states are presented in the following section through numerical analysis.
III Numerical results and discussion
@@
In this section, we numerically study the magnetic dipole moments of $\Omega_{c}$-like states. The physical parameters to be calculated are analyzed using the following parameters: $m_{u}=2.16^{+0.49}_{-0.26}$ MeV, $m_{d}=4.67^{+0.48}_{-0.17}$ MeV, $m_{s}=93.4^{+8.6}_{-3.4}\,\mbox{MeV}$, $m_{c}=1.27\pm 0.02\,$GeV [57], $\langle\bar{q}q\rangle$=$(-0.24\pm 0.01)^{3}\,$eV3, $\langle\bar{s}s\rangle$= $0.8\langle\bar{q}q\rangle$ [58], $m_{0}^{2}=0.8\pm 0.1$ GeV2, and $\langle g_{s}^{2}G^{2}\rangle=0.48\pm 0.14~{}$GeV4 [59]. The mass and residue values for these states have been calculated using the two-point sum rules of QCD, which have been used in our analysis [49, 50]. The photon distribution amplitudes and the parameters utilized in these distribution amplitudes are borrowed from Ref. [53].
@@
Beyond the parameters listed above, the QCD light-cone sum rules method includes two other parameters known as the Borel mass parameter $\rm{M^{2}}$ and the continuum threshold $\rm{s_{0}}$, and in the ideal scenario, these two parameters are expected to have a very small impact on the results. To do this, it is necessary to find the working regions of these parameters, where the variation of the results with respect to these parameters should be minimal. The $\rm{s_{0}}$ is the point at which the correlation function begins to include contributions from both excited states and the continuum. To determine the working region of this parameter, the assumption $\rm{s_{0}}=(M_{H}+0.4^{+0.1}_{-0.1})~{}\rm{GeV}^{2}$ is usually adopted. The results are then analyzed for their dependence on slight variations of this parameter. We search for the ideal Borel parameters $\rm{M^{2}}$ according to the two criteria: pole contribution (PC) and OPE convergence (CVG). These two conditions are both satisfied in the region $11.4~{}\mbox{GeV}^{2}\leq\rm{s_{0}}\leq 12.8~{}\mbox{GeV}^{2}$ and $3.0~{}\mbox{GeV}^{2}\leq\rm{M^{2}}\leq 4.2~{}\mbox{GeV}^{2}$ for the $\Omega_{c}$ states; and $11.9~{}\mbox{GeV}^{2}\leq\rm{s_{0}}\leq 13.3~{}\mbox{GeV}^{2}$ and $3.2~{}\mbox{GeV}^{2}\leq\rm{M^{2}}\leq 4.6~{}\mbox{GeV}^{2}$ for the $\Omega_{c}^{*}$ states. Our numerical evaluations show that considering these working intervals for the $\rm{M^{2}}$, the magnetic dipole moments of the $\Omega_{c}$-like states PC vary within the interval $34\%\leq\rm{PC}\leq 61\%$ and $31\%\leq\rm{PC}\leq 57\%$ for $\Omega_{c}$ and $\Omega_{c}^{*}$ states, respectively. Analyzing the CVG, one sees that the contribution of the higher twist and higher dimensional terms in the OPE is $<3\%$ of the total and the series shows good convergence for both $\Omega_{c}$ and $\Omega_{c}^{*}$ states. As an example, for fixed values of the continuum threshold $\rm{s_{0}}$, the variation of the extracted magnetic dipole moments of the spin-$\frac{1}{2}^{-}$ $\Omega_{c}$-like states with Borel mass $\rm{M^{2}}$ is shown in Fig.1. As it is seen from this figure, the magnetic dipole moments of these states exhibit a slight dependence on the parameter $\rm{M^{2}}$. While the magnetic dipole moments of the $\Omega_{c}$-like states exhibit some dependence on $\rm{s_{0}}$, this is within the limits authorized by the approach and constitutes the main part of the uncertainties.
Figure 1: Dependencies of the magnetic dipole moments of spin-$\frac{1}{2}^{-}$ $\Omega_{c}$ states on $\rm{M^{2}}$ at three values of $\rm{s_{0}}$.
@@
After setting all the relevant parameters, we are finally ready to estimate the numerical values of the corresponding magnetic dipole moments of $\Omega_{c}$-like states. The predicted results for the magnetic dipole moments are shown in Table 1. The errors of the results, caused by errors in the input parameters and uncertainties in the estimation of the working intervals of the auxiliary parameters, are also shown in Table 1.
Table 1: Results of the magnetic dipole moments of the $\Omega_{c}$-like states.
     Parameters      $J^{P}$      $ssuc\bar{u}$      $susc\bar{u}$      $ussc\bar{u}$      $ssdc\bar{d}$      $sdsc\bar{d}$      $dssc\bar{d}$
     $\mu[\mu_{N}]$      $\frac{1}{2}^{-}$      $-0.31^{+0.20}_{-0.16}$      $-1.10^{+0.23}_{-0.25}$      $-1.09^{+0.24}_{-0.23}$      $0.17^{+0.11}_{-0.08}$      $0.62^{+0.14}_{-0.14}$      $0.61^{+0.15}_{-0.13}$
     $\mu[\mu_{N}]$      $\frac{3}{2}^{-}$      $0.28^{+0.17}_{-0.13}$      $0.88^{+0.15}_{-0.18}$      $0.87^{+0.16}_{-0.16}$      $-0.058^{+0.032}_{-0.031}$      $-0.37^{+0.05}_{-0.07}$      $-0.37^{+0.05}_{-0.07}$
@@
The order of magnetic dipole moments can yield insights into their experimental measurability. The analysis of the obtained results indicates that the extracted magnetic dipole moments for both spin-1/2 and spin-3/2 $\Omega_{c}$-like states can be experimentally measured. Our analysis shows that although hadrons have the same quark content, different diquark components can significantly affect the magnetic dipole moment results. This indicates that the magnetic dipole moment is a crucial parameter in determining the internal structure of hadrons. As previously stated, the interpolating currents used for both spin-1/2 and spin-3/2 $\Omega_{c}$-like states predicted compatible with mass values. However, the magnetic dipole moments obtained for these states have different magnitudes and signs. This information may help us understand their quantum numbers and internal structure in experimental measurements. Comparing the magnetic dipole moment results obtained in this study with those extracted using different theoretical approaches can provide insight into the consistency of our predictions. The measurement of the magnetic dipole moments of unstable hadrons is a challenging experimental task. Nevertheless, using models approximating key features of QCD can give clues to the relevant dynamical mechanisms behind the observed structure. Together with spectroscopic/decay parameters and electromagnetic properties, it will also be interesting to characterize the branching ratios of the different decay modes and decay channels of $\Omega_{c}$-like states.
@@
Finally, this section briefly discusses the measurement of the magnetic dipole moment of the $\Omega_{c}$-like states in the experiment. Due to their short lifetimes, measuring the magnetic dipole moments of $\Omega_{c}$-like states by spin precession experiments is a challenging task. Instead, the magnetic dipole moment of these unstable states can only be measured indirectly through a three-step process. The particle is first produced, followed by the emission of a low-energy photon that acts as an external magnetic field. As the final step, the particle decays. There is a procedure for the indirect determination of the magnetic dipole moments of these states, which is based on the soft photon emission of hadrons proposed in Ref. [60]. The basic concept underlying the procedure is that the photon carries information about the magnetic dipole and other higher multipole moments of the emitted particle. From this point of view, we can determine the magnetic dipole moment of the particles under consideration through measuring the cross-section or decay width of the radiative process. The magnetic dipole moment of states under investigation can affect the cross sections, whether total or differential. Determining the dipole moment of states under investigation requires comparing theoretical predictions with measured cross-sections. The magnetic dipole moment of $\Delta(1232)$ resonance has been achieved by using the experimental data obtained in the $\gamma N\rightarrow\Delta(1232)\rightarrow\Delta(1232)\gamma\rightarrow\pi N\gamma$ process employing this procedure [61, 62, 63, 64, 65, 66, 67, 68, 69].

### v3

# Electromagnetic properties of $\Omega_{c}^{0}$ resonances via light-cone QCD

@@ Abstract
We systematically study the electromagnetic properties of controversial states whose internal structure is not elucidated and we try to offer a different point of view to unravel the internal structure of these states. Inspired by the $\Omega_{c}$ states observed by the LHCb Collaboration, we study the electromagnetic properties of the $\Omega_{c}$ states as the compact diquark-diquark-antiquark pentaquarks with both $J^{P}=\frac{1}{2}^{-}$ and $J^{P}=\frac{3}{2}^{-}$ in the context of the QCD light-cone sum rule model. From the obtained numerical results, we conclude that the magnetic dipole moments of the $\Omega_{c}$ states can reflect their inner structures, which can be used to distinguish their spin-parity quantum numbers. Measuring the magnetic moment of the $\Omega_{c}$ states in future experimental facilities can be very helpful for understanding the internal organization and identifying the quantum numbers of these states.
@@ I Motivation
${\Omega_{c}(3000)}:M$ $=3000.4\pm 0.2\pm 0.1\mbox{ MeV}\,,\,\,\,\,\Gamma=4.5\pm 0.6\pm 0.3\mbox{ MeV}\,,$
${\Omega_{c}(3050)}:M$ $=3050.2\pm 0.1\pm 0.1\mbox{ MeV},\,\,\,\,\,\Gamma=0.8\pm 0.2\pm 0.1\mbox{ MeV}\,,\,\,$
${\Omega_{c}(3066)}:M$ $=3065.6\pm 0.1\pm 0.3\mbox{ MeV}\,,\,\,\,\,\Gamma=3.5\pm 0.4\pm 0.2\mbox{ MeV}\,,$
${\Omega_{c}(3090)}:M$ $=3090.2\pm 0.3\pm 0.5\mbox{ MeV}\,,\,\,\,\,\Gamma=8.7\pm 1.0\pm 0.8\mbox{ MeV}\,,$
${\Omega_{c}(3119)}:M$ $=3119.1\pm 0.3\pm 0.9\mbox{ MeV}\,,\,\,\,\,\Gamma=1.1\pm 0.8\pm 0.4\mbox{ MeV}\,.$ (1)
@@
Later, the former four states $\Omega_{c}(3000)$, $\Omega_{c}(3050)$, $\Omega_{c}(3066)$ and $\Omega_{c}(3090)$ were confirmed by the LHCb and Belle Collaborations [31, 32]. The discovery of the LHCb collaboration has led to a new experimental situation, which requires a more detailed study of heavy baryons and their properties. The discovery of these states, although their quantum numbers have not been determined, could provide new insights into QCD and its complex behavior. This could lead to a deeper understanding of the underlying properties of QCD. However, our information about their properties is still not sufficient and further suggestions for experimental exploration of $\Omega_{c}$ states should be addressed. The experimental discoveries were followed by various theoretical studies, investigating them in the conventional baryon, molecular, and compact pentaquark states to shed light on their exact nature and quantum numbers (for details see the Refs. [33, 34, 35, 36, 37]).
@@
The literature review indicates that the majority of research has concentrated on computing the spectroscopic and decay parameters of these states. However, it is evident that relying exclusively on spectroscopic and decay parameters is insufficient for resolving the disputed nature of these states. Therefore, to determine the precise internal configurations of these states, additional studies such as obtaining electromagnetic multipole moments, radiative decays, and weak decays are required. To unveil the nature and internal structures of unconventional states, the physical quantities associated with the electromagnetic properties of hadrons are valuable parameters. The charge and magnetism distributions inside the hadrons can be studied using the electromagnetic form factors (FFs) and the resulting multipole moments. This information can be used to determine the distribution of quark-antiquark pairs (both valence and sea-quarks) and gluons within the hadron volume. It also provides valuable insights into the geometric forms, electric radii, and magnetic radii of the hadrons. With this motivation, in this work, we will derive the electromagnetic features of $\Omega_{c}$ states using the method of QCD light-cone sum rules [38, 39, 40]. Throughout the analysis, these states are considered in the compact $[ss][dc][\bar{d}]$, $[sd][sc][\bar{d}]$ and $[ds][sc][\bar{d}]$ diquark-diquark-antiquark configurations, and are assumed to have both $J^{P}=\frac{1}{2}^{-}$ and $J^{P}=\frac{3}{2}^{-}$ quantum numbers. There are several studies in the literature that have extracted the electromagnetic features of the singly-heavy nonconventional hadron states [41, 42, 43, 44, 45, 46, 47, 48].
@@
The article is structured as follows: we derive the QCD light-cone sum rules for the electromagnetic features of the $\Omega_{c}$ states in Sect. II. The numerical results of the magnetic dipole moments and discussions are presented in Sect. III. Sect. IV is reserved for our summary and concluding remarks.
II Formalism
@@
The study of the electromagnetic form factors of the $\Omega_{c}$ states ($\frac{1}{2}^{-}\rightarrow\Omega_{c}$, $\frac{3}{2}^{-}\rightarrow\Omega_{c}^{*}$) in QCD light cone sum rules begins with the writing of the correlation function in the weak electromagnetic background field ($\gamma$) as follows
@@
where $T$ is the time ordered product; and the $J^{\Omega_{c}}(x)$ and $J_{\mu}^{\Omega_{c}^{*}}(x)$ are the interpolating currents for the $\Omega_{c}$ states. The explicit form of the $J^{\Omega_{c}}(x)$ and $J_{\mu}^{\Omega_{c}^{*}}(x)$ are written as follows
$J^{\Omega_{c}}(x)$ $=\varepsilon^{abc}\varepsilon^{ade}\varepsilon^{bfg}\Big{\{}\big{[}q^{d^{T}}_{1}(x)C\gamma_{\mu}q_{2}^{e}(x)\big{]}\big{[}q_{3}^{f^{T}}(x)C\gamma^{\mu}c^{g}(x)\big{]}C\bar{d}^{c^{T}}(x)\Big{\}},$ (4)
$J_{\mu}^{\Omega_{c}^{*}}(x)$ $=\varepsilon^{abc}\varepsilon^{ade}\varepsilon^{bfg}\Big{\{}\big{[}q^{d^{T}}_{1}(x)C\gamma_{\mu}q_{2}^{e}(x)\big{]}\big{[}q_{3}^{f^{T}}(x)C\gamma_{5}c^{g}(x)\big{]}C\bar{d}^{c^{T}}(x)\Big{\}},$ (5)
where the $C$ is the charge conjugation operator and $a$, $b$… are color indices. The quark content of the $\Omega_{c}$ states is presented in Table 1. It should be noted here that the mass values obtained with these interpolating currents are compatible with the mass values of the $\Omega_{c}(3050)$, $\Omega_{c}(3066)$, $\Omega_{c}(3090)$, and $\Omega_{c}(3119)$ states [49, 50].
Table 1: The quark content of the $\Omega_{c}$ states.
     States      $q_{1}$      $q_{2}$      $q_{3}$
     $\big{[}ss\big{]}\big{[}dc\big{]}\bar{d}$      s      s      d
     $\big{[}sd\big{]}\big{[}sc\big{]}\bar{d}$      s      d      s
     $\big{[}ds\big{]}\big{[}sc\big{]}\bar{d}$      d      s      s
@@
In order to express the hadronic representation in a more clear and straightforward manner, it is possible to present the above equations as follows:
$\Pi^{Had}(p,q)$ $=\frac{\lambda^{2}_{\mathrm{\Omega_{c}}}}{[(p+q)^{2}-m^{2}_{\mathrm{\Omega_{c}}}][p^{2}-m^{2}_{\mathrm{\Omega_{c}}}]}\bigg{[}\Big{(}f_{1}(q^{2})+f_{2}(q^{2})\Big{)}p\!\!\!/\varepsilon\!\!\!/q\!\!\!/+\cdots\bigg{]},$ (16)
$\Pi^{Had}_{\mu\nu}(p,q)$ $=\frac{\lambda_{{}_{{\mathrm{\Omega_{c}^{*}}}}}^{2}}{[(p+q)^{2}-m_{{}_{{\mathrm{\Omega_{c}^{*}}}}}^{2}][p^{2}-m_{{}_{{\mathrm{\Omega_{c}^{*}}}}}^{2}]}\bigg{[}\varepsilon_{\mu}q_{\nu}q\!\!\!/\,F_{1}(q^{2})+g_{\mu\nu}p\!\!\!/\varepsilon\!\!\!/q\!\!\!/\,F_{2}(q^{2})+\cdots\bigg{]}.$ (17)
@@
where $F_{M}(0)$ and $G_{M}(0)$ stand for magnetic form factor of the spin-$\frac{1}{2}$ and spin-$\frac{3}{2}$ $\Omega_{c}$ states, respectively. Explicit forms of these terms are given as,
@@
$+\rm{Tr}\Big{[}\gamma^{\mu}S_{c}^{gg^{\prime}}(x)\gamma^{\nu}\widetilde{S}_{q_{2}q_{3}}^{ef^{\prime}}(x)\gamma_{\mu}S_{q_{1}q_{2}}^{de^{\prime}}(x)\gamma_{\mu}\widetilde{S}_{q_{3}q_{1}}^{fd^{\prime}}(x)\Big{]}\bigg{\}}\widetilde{S}_{d}^{c^{\prime}c}(-x)|0\rangle_{\gamma},$ (22)
@@
$+\rm{Tr}\Big{[}\gamma_{5}S_{c}^{gg^{\prime}}(x)\gamma_{5}\widetilde{S}_{q_{2}q_{3}}^{ef^{\prime}}(x)\gamma_{\mu}S_{q_{1}q_{2}}^{de^{\prime}}(x)\gamma_{\mu}\widetilde{S}_{q_{3}q_{1}}^{fd^{\prime}}(x)\Big{]}\bigg{\}}\widetilde{S}_{d}^{c^{\prime}c}(-x)|0\rangle_{\gamma},$ (23)
@@
where one of the propagators for either the light or heavy quark interacts with the photon at a short distance, and the surviving four propagators are assumed to be free. In the case of the non-perturbative contributions, the following equation is to be used
@@
After completing the procedures described above, we have finished analytical calculations of the magnetic dipole moments for the relevant states. For simplicity, analytical calculations for the spin-1/2 $\Omega_{c}$ states are presented as an example in the appendix. The results for the magnetic dipole moments of $\Omega_{c}$ states are presented in the following section through numerical analysis.
III Numerical results
@@
In this section, we numerically study the magnetic dipole moments of $\Omega_{c}$ states. The physical parameters to be calculated are analyzed using the following parameters: $m_{s}=93.4^{+8.6}_{-3.4}\,\mbox{MeV}$, $m_{c}=1.27\pm 0.02\,$GeV [57], $\langle\bar{q}q\rangle$=$(-0.24\pm 0.01)^{3}\,$eV3, $\langle\bar{s}s\rangle$= $0.8\langle\bar{q}q\rangle$ [58], $m_{0}^{2}=0.8\pm 0.1$ GeV2, and $\langle g_{s}^{2}G^{2}\rangle=0.48\pm 0.14~{}$GeV4 [59]. In numerical analysis, we set $m_{u}$ =$m_{d}$ = 0 and $m^{2}_{s}=0$, but consider terms proportional to $m_{s}$. The mass and residue values for these states have been calculated using the two-point sum rules of QCD, which have been used in our analysis [49, 50]. The photon distribution amplitudes and the parameters utilized in these distribution amplitudes are borrowed from Ref. [53].
@@
Beyond the parameters listed above, the QCD light-cone sum rules method includes two other parameters known as the Borel mass parameter $\rm{M^{2}}$ and the continuum threshold $\rm{s_{0}}$, and in the ideal scenario, these two parameters are expected to have a very small impact on the results. To do this, it is necessary to find the working regions of these parameters, where the variation of the results with respect to these parameters should be minimal. The $\rm{s_{0}}$ is the point at which the correlation function begins to include contributions from both excited states and the continuum. To determine the working region of this parameter, the assumption $\rm{s_{0}}=(M_{H}+0.4^{+0.1}_{-0.1})~{}\rm{GeV}^{2}$ is usually adopted. The results are then analyzed for their dependence on slight variations of this parameter. We search for the ideal Borel parameters $\rm{M^{2}}$ according to the two criteria: pole contribution (PC) and OPE convergence (CVG). These two conditions are both satisfied in the region $11.4~{}\mbox{GeV}^{2}\leq\rm{s_{0}}\leq 12.8~{}\mbox{GeV}^{2}$ and $3.0~{}\mbox{GeV}^{2}\leq\rm{M^{2}}\leq 4.2~{}\mbox{GeV}^{2}$ for the $\Omega_{c}$ states; and $11.9~{}\mbox{GeV}^{2}\leq\rm{s_{0}}\leq 13.3~{}\mbox{GeV}^{2}$ and $3.2~{}\mbox{GeV}^{2}\leq\rm{M^{2}}\leq 4.8~{}\mbox{GeV}^{2}$ for the $\Omega_{c}^{*}$ states. Our numerical evaluations show that considering these working intervals for the $\rm{M^{2}}$, the magnetic dipole moments of the $\Omega_{c}$ states PC vary within the interval $34\%\leq\rm{PC}\leq 61\%$ and $31\%\leq\rm{PC}\leq 57\%$ for $\Omega_{c}$ and $\Omega_{c}^{*}$ states, respectively. Analyzing the CVG, one sees that the contribution of the higher twist and higher dimensional terms in the OPE is $<3\%$ of the total and the series shows good convergence for both $\Omega_{c}$ and $\Omega_{c}^{*}$ states. As an example, for fixed values of the continuum threshold $\rm{s_{0}}$, the variation of the extracted magnetic dipole moments of the spin-$\frac{1}{2}^{-}$ $\Omega_{c}$ states with Borel mass $\rm{M^{2}}$ is shown in Fig.1. As it is seen from this figure, the magnetic dipole moments of these states exhibit a slight dependence on the parameter $\rm{M^{2}}$. While the magnetic dipole moments of the $\Omega_{c}$ states exhibit some dependence on $\rm{s_{0}}$, this is within the limits authorized by the approach and constitutes the main part of the uncertainties.
@@
After setting all the relevant parameters, we are finally ready to estimate the numerical values of the corresponding magnetic dipole moments of $\Omega_{c}$ states. The predicted results for the magnetic dipole moments are shown in Table 2. The errors of the results, caused by errors in the input parameters and uncertainties in the estimation of the working intervals of the auxiliary parameters, are also shown in Table 2.
Table 2: Results of the magnetic dipole moments of the $\Omega_{c}$ states.
     Parameter      $J^{P}$      $ssdc\bar{d}$      $sdsc\bar{d}$      $dssc\bar{d}$
     $\mu[\mu_{N}]$      $\frac{1}{2}^{-}$      $0.17^{+0.11}_{-0.08}$      $0.62^{+0.14}_{-0.14}$      $0.61^{+0.15}_{-0.13}$
     $\mu[\mu_{N}]$      $\frac{3}{2}^{-}$      $-0.058^{+0.032}_{-0.031}$      $-0.37^{+0.05}_{-0.07}$      $-0.37^{+0.05}_{-0.07}$
@@
The order of magnetic dipole moments can yield insights into their experimental measurability. The analysis of the obtained results indicates that the extracted magnetic dipole moments for both spin-1/2 and spin-3/2 $\Omega_{c}$ states can be experimentally measured. Our analysis shows that although hadrons have the same quark content, different diquark components can significantly affect the magnetic dipole moment results. This indicates that the magnetic dipole moment is a crucial parameter in determining the internal structure of hadrons. As previously stated, the interpolating currents used for both spin-1/2 and spin-3/2 $\Omega_{c}$ states predicted compatible with mass values. However, the magnetic dipole moments obtained for these states have different magnitudes and signs. This information may help us understand their quantum numbers and internal structure in experimental measurements. Comparing the magnetic dipole moment results obtained in this study with those extracted using different theoretical approaches can provide insight into the consistency of our predictions. The measurement of the magnetic dipole moments of unstable hadrons is a challenging experimental task. Nevertheless, using models approximating key features of QCD can give clues to the relevant dynamical mechanisms behind the observed structure. Together with spectroscopic/decay parameters and electromagnetic properties, it will also be interesting to characterize the branching ratios of the different decay modes and decay channels of $\Omega_{c}$ states.
@@
Finally, this section briefly discusses the measurement of the magnetic dipole moment of the $\Omega_{c}$ states in the experiment. Due to their short lifetimes, measuring the magnetic dipole moments of $\Omega_{c}$ states by spin precession experiments is a challenging task. Instead, the magnetic dipole moment of these unstable states can only be measured indirectly through a three-step process. The particle is first produced, followed by the emission of a low-energy photon that acts as an external magnetic field. As the final step, the particle decays. There is a procedure for the indirect determination of the magnetic dipole moments of these states, which is based on the soft photon emission of hadrons proposed in Ref. [60]. The basic concept underlying the procedure is that the photon carries information about the magnetic dipole and other higher multipole moments of the emitted particle. From this point of view, we can determine the magnetic dipole moment of the particles under consideration through measuring the cross-section or decay width of the radiative process. The magnetic dipole moment of states under investigation can affect the cross sections, whether total or differential. Determining the dipole moment of states under investigation requires comparing theoretical predictions with measured cross-sections. The magnetic dipole moment of $\Delta(1232)$ resonance has been achieved by using the experimental data obtained in the $\gamma N\rightarrow\Delta(1232)\rightarrow\Delta(1232)\gamma\rightarrow\pi N\gamma$ process employing this procedure [61, 62, 63, 64, 65, 66, 67, 68].